OpenStudy (anonymous):

Solve by completing the square: 3 + 2x^2 = -8x

OpenStudy (anonymous):

probably want to star with \[2x^2+8x=-3\] then divide by \(2\) and get \[x^2+4x=-\frac{3}{2}\]

OpenStudy (anonymous):

then complete the square as usual do you know how to do that?

OpenStudy (anonymous):

Ok, and no i dont

OpenStudy (anonymous):

ok then lets do it what is half of 4?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

ok then write \[(x+2)^2=-\frac{3}{2}+?\] now what is 2 squared

OpenStudy (anonymous):

4

OpenStudy (anonymous):

yes, so write \[(x+2)^2=-\frac{3}{2}+4=\frac{5}{2}\] take the square root of both sides get \[x+2=\pm\sqrt{\frac{5}{2}}\] then subtract \(2\) from both sides and you are done

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

yw